2018-05-14-CSw » History » Revision 3

« Previous | Revision 3/4 (diff) | Next »
Rajaram, Durga, 15 May 2018 15:24

Computing and Software status, May 15, 2018

  • Software:
    • TOF:
      • one PMT channel with hole in the ADC distribution -- needs to be blacklisted & accounted for in the reconstruction
      • Calibrations need to be improved & finalized to try improve efficiency especially for TOF2
        • VP having some issues, DR to interface with him
      • TOF review has been on hiatus, to resume
  • Tracker:
    • Geometry description issue
    • Some evidence (?) that the tracker is misaligned for the scattering run period
  • EMR
    • bug fix: chi2 calculation
    • bug fix: polystyrene material definition
  • Geometry:
    • LH2 window material is incorrect in production geometries -- fix in hand
    • Tracker glue not correctly defined in tracker geometry -- CR has tuned the density
    • DR to push fixes to all Step IV production geometries
    • Standing: description of diffuser aperture & support materials
  • MC
    • bug fix: we were not seeing delta electrons due to a bug -- fixed in MAUS trunk
    • bug fix: tracker channel/calibration look up was incorrect in MC digitizer -- fixed in trunk
  • Processing:
    • Plan:
      • make a new release
      • push corrected geometries
      • run MC production
      • reprocess data
  • Computing:
    • MLCR decomissioning (PF)
      • Equipment & services to be migrated
        • Target: basically done (tested and training to be finalized)
        • CDB: failover next week. One more week for the PPD slave in order to be independent from the MLCR.
        • Backups on GRID: tarballs ready for the Archivist to be moved
        • Switch off: D1 survey next week before turning most of the machines off and final Archiver tarball. PF waiting for survey before turning off IOCs and machines.
      • Computing equipment to be returned
        • QPS computer
        • Geneva computers -- onrec machines being used for local processing, so we'd like to hold on to them as long as possible
  • Staffing
    • HN's grid roles: agreed that PK + student/Brunel person will take over from Henry
    • DM says he will be able to continue supporting GRID productions/processing & if support needed he can lean on Mihailo
  • MAUS Paper:
    • (Finally) revised draft sent to KL, PF

Updated by Rajaram, Durga about 6 years ago · 3 revisions